Suppose the Blue Hand Hospitality Enterprises Co. wishes to set aside an equal, annual, end-of-year amount in an investment account earning 9.50% per year over the next five years. The firm wants to have $5 million in the account when the account matures. How much must be deposited in the account at the end of each year?

Group of answer choices

$827,182.09

$891,410.73

$794,569.44

$844,200.00
Answer #1

- Future Value required in 20 years is \(\$ 5,000,000\)

Calculating the Annual Deposit to be made each year using Fv of ordinary annuity formula:-

FutureValue \(=C * \frac{\left[(1+r)^{n}-1\right]}{r}\)

Where, \(\mathrm{C}=\) Periodic Deposit

\(r=\) Periodic Interest rate \(=9.50 \%\)

\(n=\) no of periods \(=5\)

\(5,000,000=C * \frac{\left[(1+0.095)^{5}-1\right]}{0.095}\)

\(5,000,000=C * \frac{[1.57423874093-1]}{0.095}\)

\(C=\$ 827,182.09\)

So, deposited in the account at the end of each year is \(\$ 827,182.09\)

Option 1

If you need any clarification, you can ask in comments.